Factor out the GCF in each expression.
(x - 2)x -(x - 2)

Factor out the GCF in each expression, then factor out the opposite of the GCF.
12u^5v^6 + 18u^2^v3 - 15u^4v^5

Answer by user_dude2008(1825)

(x - 2)x -(x - 2) = (x-2)(x-1)



12u^5v^6 + 18u^2^v3 - 15u^4v^5 = 3u^2v^3(4u^3v^3+6-5u^2v^2)

12u^5v^6 + 18u^2^v3 - 15u^4v^5 = -3u^2v^3(-4u^3v^3-6+5u^2v^2)